# Classically-embedded split Cayley hexagons rule three-qubit contextuality with three-element contexts

### v1

@@ 4.4 Classically-embedded hexagons encode three-qubit
Let start with elliptic quadrics. By a computer search we have found that each elliptic quadric features 9 pairwise disjoint unsatisfiable lines/contexts forming a spread. On the other hand, each classically embedded hexagon shares with each elliptic quadric such a set of 9 lines. An example is illustrated in Figure 11.
@@
Figure 11: The generic intersection of a classical hexagon with an elliptic quadric: nine mutually disjoint lines that are at maximum distance from each other, aka a distance-three spread of the hexagon. Note that these nine lines cover all the 27 points of the quadric.
@@
Next, each classically embedded hexagon shares with each hyperbolic quadric exactly 21 lines, forming a pattern isomorphic to the one shown in Figure 12. People familiar with graph theory may readily recognize this pattern as the Heawood graph, also known as the point-line incidence graph of the Fano plane, whose vertices are represented by bigger bullets, each of its edges being supplied with one more point/observable to represent a full line of the three-qubit $\mathcal{W}(5,2)$. Using a computer we have verified that each three-qubit hyperbolic quadric features 21 unsatisfiable lines whose arrangement is isomorphic to that depicted in Figure 12.
@@
Figure 12: The generic intersection of a classical hexagon with a hyperbolic quadric: 21 lines forming edges of the Heawood graph. As in the previous case, the 21 lines comprise all the 35 points of the quadric.
@@
A similar analysis carried out for doilies living in $\mathcal{W}(5,2)$ showed the same result. The three unsatisfiable constraints a doily features, forming always a spread in one of the doily’s grids, are those three lines that the doily shares with a particular copy of a classical hexagon.
@@
The message from the above-given observations is more than obvious: the degree of contextuality of a contextual quantum geometry of the three-qubit $\mathcal{W}(5,2)$ whose contexts are i) all the 315 lines of the space, ii) all the lines of an elliptic quadric, iii) all the lines of a hyperbolic quadric or iv) all the lines of a doily is equal to the number of lines each of these geometries shares with a classically embedded hexagon, with the understanding that each shared line corresponds to an unsatisfiable context. A natural generalization of these findings is that this should hold for any contextual sub-geometry of $\mathcal{W}(5,2)$ whose contexts are lines. In other words, classically-embedded copies of $\mathcal{H}$, although being non-contextual by themselves, are found to rule three-qubit contextuality by being a reliable tool not only to check whether a particular subgeometry of $\mathcal{W}(5,2)$ is contextual, but also, and still mysteriously, to ‘extract’ from this subgeometry exactly that part that makes it non-contextual!
@@
As a particular stance of this conjecture, let us consider the complement of a skew-embedded hexagon, which is a contextual $(63_{12},252_{3})$-configuration [holweck2022three]. Analyzing complements of several different copies of skew-embedded hexagons we always arrived at the same result: each complement featured 24 unsatisfiable constraints that were exactly those 24 lines in which the given skew-embedded hexagon differs from the classical one derived from it using the procedure described in Section 4.3; thus, for example, if we consider the complement of the skew hexagon shown in Figure 9 then its 24 unsatisfiable lines are exactly the 24 gray lines of its derived classical sibling portrayed in Figure 10!
5 Some experimental verification using the
IBM Quantum Experience
@@ 5 Some experimental verification using the
The advent of accessible NISQ computers has allowed researchers to test whether contextual geometries exhibit the predicted quantum behaviour in reality, see for example [laghaout2022demonstration, holweck2021testing, Kelleher_2_qubit_games]. Given the accelerated development of such technologies, and their widening availability for research purposes, it is easier than ever to experimentally test out the predictions of NCHV models mentioned in the introduction. First, one needs a quantitative test on our geometry that can rule out NCVM models in favour of those predicted by quantum mechanics (QM). In this section we examine such a test.
@@ 5.1 Cabello inequality
In 2010 Cabello [Cab10] introduced an inequality dependent upon the number of satisfiable constraints in a contextual geometry. The quantity $\chi$ is defined as the sum of expectation values of all constraints, with negative ones picking up a sign change,
@@
The inequalities provide an upper bound on the measured value of $\chi$ based on whether all constraints are satisfiable (QM) or there is a NCHV model constraining some measurement outcomes (HV).
5.2 Measuring contextuality on a NISQ computer
One can experimentally test the contextuality (that is, the value of $\chi$ in (16)) of a given labelled geometry via the IBM Quantum Experience [ibmq]. The methodology is straightforward: for a given context in a geometry labelled by $3$-qubit operators, measure the values of the operators at each point of the context and combine to get the measured parity of the line. Repeat for all lines in the geometry, and sum together weighted by their signed parities. If the final result is greater than the upper bound $N-2d$ given in (16), then we have demonstrated the contextuality of the geometry.
@@ 5.2 Measuring contextuality on a NISQ computer
Note that in both cases our experiments reveal the contextual nature of the configuration. Our procedure to test Cabello’s inequalities follow and slightly improves the experiments of [laghaout2022demonstration, holweck2021testing]. In particular, this time we delegated the measurements corresponding to the three observables of a context to three different delegation qubits (instead of a single qubit used in the previous cited work). Although this new procedure reduces the gap between the experimental value and the classical bound, it has the advantage of collecting the result for each of the three observables independently.
@@ 6 Conclusion
This paper provides substantial insights into the role played by the smallest split Cayley hexagon $\mathcal{H}$ in its classical embeddings into $\mathcal{W}(5,2)$ in observable-based three-qubit contextuality, as first elaborated by three of us in [holweck2022three]. We have demonstrated that classically-embedded copies of $\mathcal{H}$ fully encode the basic information about distinguished contextual configurations living in $\mathcal{W}(5,2)$, namely doilies, both types of quadrics, complements of skew-embedded $\mathcal{H}$’s as well as the configuration formed by all the 315 lines of the space. Given such a configuration, one can always find some classically-embedded copy of $\mathcal{H}$ that shares with this configuration exactly those contexts that are unsatisfiable by an NCHV model! It is truly amazing to realize that the three-qubit $\mathcal{W}(5,2)$ is endowed with the distinguished subgeometry, $\mathcal{H}$, that – although being itself non-contextual – is able to single out from a large variety of contextual configurations exactly those parts that ‘responsible’ for their contextual behavior.
@@
Interestingly, we have already at hand some hints that something similar is taking place in the next case in the hierarchy – the four-qubit $\mathcal{W}(7,2)$. By making use of the Lagrangian Grassmannian mapping of type $LGr(3,6)$ that sends planes of $\mathcal{W}(5,2)$ into points of a hyperbolic quadric $\mathcal{Q}^{+}(7,2)$ of $\mathcal{W}(7,2)$ we already found on this quadric a particular configuration that can be regarded as a four-qubit analog of a three-qubit Heawood-graph-underpinned configuration described in Section 4.4 (cf. Figure 12). This particular configuration contains 135 points and 315 lines, with seven lines through a point and three points on a line, and – being isomorphic to the dual polar space $\mathcal{DW}(5,2)$ – has the desired property that it shares with each of the 120 $\mathcal{W}(5,2)$’s located on the $\mathcal{Q}^{+}(7,2)$ a copy of $\mathcal{H}$, the latter being indeed embedded classically into the corresponding $\mathcal{W}(5,2)$; moreover, it also contains 36 (one per each hyperbolic quadric of $\mathcal{W}(5,2)$ as dictated by $LGr(3,6)$-correspondence) copies of the point-plane incidence graph of PG$(3,2)$. What remains to be checked is whether there exists a four-qubit analog of a ‘classical’ $\mathcal{H}$ of $\mathcal{W}(5,2)$, that is a configuration that picks up from each $\mathcal{Q}^{+}(7,2)$ of $\mathcal{W}(7,2)$ a configuration isomorphic to our $(135_{7},315_{3})$-one. An affirmative answer to this computationally rather challenging task would mean, among other things, that the degree of contextuality of a four-qubit hyperbolic quadric is 315 and that of the whole $\mathcal{W}(7,2)$ amounts to 1 575.
@@ Acknowledgments
This work was supported in part by the Slovak VEGA Grant Agency, project number 2/0004/20, by the PEPR integrated project EPiQ ANR-22-PETQ-0007 part of Plan France 2030 and by the project TACTICQ of the EIPHI Graduate School (contract ANR-17-EURE-0002). We acknowledge the use of the IBM Quantum Experience and the IBMQ-research program. The views expressed are those of the authors and do not reflect the official policy or position of IBM or the IBM Quantum Experience team. One would like to thank the developers of the open-source framework Qiskit. We also thank Dr. Petr Pracna (Technology Centre ASCR, Prague, Czech Republic) for his invaluable help in preparing Figures 5 to 12 as well as Dr. Péter Lévay (Budapest University of Technology, Budapest, Hungary) for sharing his view on the action of $Sp(6,2)$ on $\mathcal{W}(5,2)$.

### v3

@@ 4.4 Classically-embedded hexagons encode three-qubit contextuality
Let start with elliptic quadrics. By a computer search we have found that each elliptic quadric features 9 pairwise disjoint unsatisfiable lines/contexts forming a spread. On the other hand, each classically-embedded hexagon shares with each elliptic quadric such a set of 9 lines. An example is illustrated in Figure 12.
@@
Figure 12: The generic intersection (bold-faced points and lines) of a classical hexagon with the elliptic quadric whose index is $YYY$: nine mutually disjoint lines that are at maximum distance from each other. Note that these nine lines cover all the 27 points of the quadric.
@@
Next, each classically-embedded hexagon shares with each hyperbolic quadric exactly 21 lines, forming a pattern isomorphic to the one shown in Figure 13. By comparing this figure with Figure 5 one readily recognizes this pattern as the Heawood graph, also known as the point-line incidence graph of the Fano plane, whose vertices are represented by bigger bullets, each of its edges being supplied with one more point/observable to represent a full line of the three-qubit $\mathcal{W}(5,2)$. Using a computer-aided search based on a SAT solver we have verified that each three-qubit hyperbolic quadric features 21 unsatisfiable lines whose arrangement is isomorphic to that depicted in Figure 13.
@@
Figure 13: The generic intersection (bold-faced points and lines) of a classical hexagon with the hyperbolic quadric whose index is $III$: 21 lines forming edges of the Heawood graph (compare also with Figure 5). As in the previous case, the 21 lines comprise all the 35 points of the quadric.
@@
A similar analysis carried out for doilies living in $\mathcal{W}(5,2)$ showed the same result. The three unsatisfiable constraints a doily features are those three lines that the doily shares with a particular copy of the classical hexagon. An apt example of this property is furnished by the quadratic doily that is shared by the elliptic quadric of index $YYY$ and the hyperbolic quadric of index $III$. By comparing Figure 12 and Figure 13 one can see that there are indeed just three boldfaced lines that occur in both the figures, namely the lines $IXX-IYY-IZZ$, $XIX-YIY-ZIZ$ and $ZZI-YYI-XXI$.
@@
The message from the above-given observations is more than obvious: the degree of contextuality of a contextual quantum geometry of the three-qubit $\mathcal{W}(5,2)$ whose contexts are i) all the 315 lines of the whole space $\mathcal{W}(5,2)$, ii) all the 45 lines of an elliptic quadric, iii) all the 105 lines of a hyperbolic quadric or iv) all the 15 lines of a doily is equal to the number of lines each of these geometries shares with a classically-embedded hexagon, with the understanding that each shared line corresponds to an unsatisfiable context. A natural generalization of these findings is that this should hold for any contextual sub-geometry of $\mathcal{W}(5,2)$ whose contexts are lines. In other words, classically-embedded copies of $\mathcal{H}$, although being non-contextual by themselves, are found to rule three-qubit contextuality with three-element contexts by being a reliable tool not only to check whether a particular subgeometry of $\mathcal{W}(5,2)$ is contextual, but also, and still mysteriously, to ‘extract’ from this subgeometry exactly that part that makes it non-contextual!
@@
As a particular stance of this conjecture, let us consider the complement of a skew-embedded hexagon, which is a contextual $(63_{12},252_{3})$-configuration [18]. Analyzing complements of several different copies of skew-embedded hexagons we always arrived at the same result: each complement featured 24 unsatisfiable constraints that were exactly those 24 lines in which the given skew-embedded hexagon differs from the classical one derived from it using the procedure described in Section 4.3; thus, for example, if we consider the complement of the skew hexagon shown in Figure 10 then its 24 unsatisfiable lines are exactly the 24 gray lines of its derived classical sibling portrayed in Figure 11!
5 Some experimental verification using the IBM Quantum Experience
@@
The advent of accessible NISQ computers has allowed researchers to test whether contextual geometries exhibit the predicted quantum behaviour in reality, see for example [22, 17, 20, 42, 5, 2, 48]. Given the accelerated development of such technologies, and their widening availability for research purposes, it is easier than ever to experimentally test out the predictions of NCHV models mentioned in the introduction. First, one needs a quantitative test on our geometry that can rule out NCHV models in favour of those predicted by quantum mechanics (QM). In this section we examine such a test.
@@ 5.1 Cabello inequality
In 2010 Cabello [7] introduced an inequality dependent upon the number of satisfiable constraints in a contextual geometry. The quantity $\chi$ is defined as the sum of expectation values of all constraints, with negative ones picking up a sign change,
@@
The inequalities provide an upper bound on the measured value of $\chi$ based on whether all constraints are satisfiable (QM) or there is an NCHV model constraining some measurement outcomes (HV).
5.2 Measuring contextuality on an NISQ computer
One can experimentally test the contextuality (that is, the value of $\chi$ in (21)) of a given labelled geometry via the IBM Quantum Experience [1]. The methodology is straightforward: for a given context in a geometry labelled by $3$-qubit operators, measure the values of the operators at each point of the context and combine to get the measured parity of the line. Repeat for all lines in the geometry, and sum together weighted by their signed parities. If the final result is greater than the upper bound $N-2d$ given in (21), then we have demonstrated the contextuality of the geometry.
@@ 5.2 Measuring contextuality on an NISQ computer
Note that in both cases our experiments reveal the contextual nature of the configuration. Our procedure to test Cabello’s inequalities follow and slightly improves the experiments of [22, 17]. In particular, this time we delegated the measurements corresponding to the three observables of a context to three different delegation qubits (instead of a single qubit used in the previous cited work). Although this new procedure reduces the gap between the experimental value and the classical bound, it has the advantage of collecting the result for each of the three observables independently.
@@ 6 Conclusion
This paper provides substantial insights into the role played by the smallest split Cayley hexagon $\mathcal{H}$ in its classical embeddings into $\mathcal{W}(5,2)$ in observable-based three-qubit contextuality, as first elaborated by three of us in [18]. We have demonstrated that classically-embedded copies of $\mathcal{H}$ fully encode the basic information about distinguished contextual configurations living in $\mathcal{W}(5,2)$, namely doilies, both types of quadrics, complements of skew-embedded $\mathcal{H}$’s as well as the configuration formed by all the 315 lines of the space. Given such a configuration, one can always find some classically-embedded copy of $\mathcal{H}$ that shares with this configuration exactly those contexts that are unsatisfiable by an NCHV model! It is truly amazing to realize that the three-qubit $\mathcal{W}(5,2)$ is endowed with the distinguished subgeometry, $\mathcal{H}$, that – although being itself non-contextual – is able to single out from a large variety of contextual configurations exactly those parts that ‘responsible’ for their contextual behavior.
@@
Interestingly, we have already at hand some hints that something similar is taking place in the next case in the hierarchy – the four-qubit $\mathcal{W}(7,2)$ [30]. By making use of the Lagrangian Grassmannian mapping of type $LGr(3,6)$ that sends planes of $\mathcal{W}(5,2)$ into points of a hyperbolic quadric $\mathcal{Q}^{+}(7,2)$ of $\mathcal{W}(7,2)$ we already found on this quadric a particular configuration that can be regarded as a four-qubit analog of a three-qubit Heawood-graph-underpinned configuration described in Section 4.4 (cf. Figure 13). This particular configuration contains 135 points and 315 lines, with seven lines through a point and three points on a line, and – being isomorphic to the dual polar space $\mathcal{DW}(5,2)$ – has the desired property that it shares with each of the 120 $\mathcal{W}(5,2)$’s located on the $\mathcal{Q}^{+}(7,2)$ a copy of $\mathcal{H}$, the latter being indeed embedded classically into the corresponding $\mathcal{W}(5,2)$; moreover, it also contains 36 (one per each hyperbolic quadric of $\mathcal{W}(5,2)$ as dictated by $LGr(3,6)$-correspondence) copies of the point-plane incidence graph of PG$(3,2)$. What remains to be checked is whether there exists a four-qubit analog of a ‘classical’ $\mathcal{H}$ of $\mathcal{W}(5,2)$, that is a configuration that picks up from each $\mathcal{Q}^{+}(7,2)$ of $\mathcal{W}(7,2)$ a configuration isomorphic to our $(135_{7},315_{3})$-one. An affirmative answer to this computationally rather challenging task would mean, among other things, that the degree of contextuality of a four-qubit hyperbolic quadric is 315 and that of the whole $\mathcal{W}(7,2)$ amounts to 1 575.
Finally, as suggested by one of the reviewers, it would be worth having a closer look if there is something deeper behind a seemingly formal analogy between the two inequivalent embeddings of the smallest split Cayley hexagon into the three-qubit $\mathcal{W}(5,2)$ and two inequivalent kinds of genuine tripartite entanglement, represented by the $GHZ$ state and the $W$ state [11, 8]. As it is well known, the entanglement of the $GHZ$ state [13] disappears if any of the three qubits is traced over, whereas the entanglement in the $W$ state survives the loss of any of the three qubits; in other words, the entanglement in the $W$ state is robust against the loss of one qubit, while the $GHZ$ state is reduced to a product of two qubits. Let us perform a similar trace-out-a-single-qubit procedure on the 63 observables of a three-qubit hexagon and see which out of the 63 lines of such a hexagon retain their totally-isotropic character, i. e. reduce to two-qubit lines. A brief inspection shows that in the case of a classically-embedded hexagon these two-qubit lines cover all the points except for those particular three ones that feature $I$ on both non-traced-out positions. A skew-embedded hexagon contains, however, additional points that are left out, forming several different patterns. One of them features six additional points that lie in triples on two disjoint lines; moreover, and quite remarkably, these two lines define a unique grid whose third line that is disjoint from the two is nothing but the axis of the corresponding hexagon. So, like in the entanglement case, this observable-related tracing-out procedure leads to different outcomes in dependence on the way a hexagon is embedded into $\mathcal{W}(5,2)$. The very fact that the axis of a skew-embedded hexagon emerges also in this analogy is of particular interest. For example, it may lend itself as a sort of guiding principle when addressing the above-discussed four-qubit case: one can assume that some of the embeddings of a yet-to-be-discovered geometric configuration ruling four-qubit contextuality could well feature a distinguished linear subspace having one more dimension – i. e. a plane. Moreover, given the fact that there are as many as nine inequivalent forms of four-qubit entanglement [44], our ruling configuration should have eight more companions having the same underlying geometry but being embedded differently into $\mathcal{W}(7,2)$. And this is certainly a topic that deserves to be properly treated in a separate paper.
@@ Acknowledgments
This work was supported in part by the Slovak VEGA Grant Agency, project number 2/0043/24, by the PEPR integrated project EPiQ ANR-22-PETQ-0007 part of Plan France 2030 and by the project TACTICQ of the EIPHI Graduate School (contract ANR-17-EURE-0002). We acknowledge the use of the IBM Quantum Experience and the IBMQ-research program. The views expressed are those of the authors and do not reflect the official policy or position of IBM or the IBM Quantum Experience team. One would like to thank the developers of the open-source framework Qiskit. We also thank Mr. Zsolt Szabó (Robert Bosch Kft, Budapest, Hungary) for the help with Figure 5, Dr. Petr Pracna (Technology Centre ASCR, Prague, Czech Republic) for his invaluable help in preparing Figures 6 to 13 as well as Dr. Péter Lévay (Budapest University of Technology, Budapest, Hungary) for sharing his view on the action of $Sp(6,2)$ on $\mathcal{W}(5,2)$. Last but not least, we are also grateful to the reviewers for their useful remarks/suggestions.
Data availability
All codes and results of our quantum experiments are openly available at: https://quantcert.github.io/hexagon. The computations of degrees of contextuality are done with the software Qontextium, openly available at https://quantcert.github.io/contextualityDegree/.